Determine the empirical and molecular formula of a compound composed of 18.24 g Carbon, 0.51 g Hydrogen, and 16.91 g Fluorine has a molar mass 562.0 g/mol.

1
Expert's answer
2021-11-19T16:01:01-0500

Element

  1. Carbon
  2. Hydrogen
  3. Flourine


Mass

  1. 18.24
  2. 0.51
  3. 16.91


Molecular mass

  1. 12
  2. 1
  3. 18.998


Moles 

  1. (18.24÷12)=1.52
  2. (0.51÷1)=0.51
  3. (16.91 ÷18.998) =0.890


Mole ratio

  1. (1.52÷0.51)=3
  2. (0.51÷0.51)=1
  3. (0.89 ÷0.51) =1.75 =7/4


C=3 H=1 F=7/4

C=12 H=4 F=7

Empirical formula =C12H4H7

Empirical formula mass=(12*12)+ (4*1) + (7*19)=281

Mass formula=562

Therefore, mass formula /empirical formula=562÷281=2

Compound formula =C24H8F14